HUBBLE SPACE TELESCOPE DAILY REPORT #2040 PERIOD COVERED: 0000Z (UTC) 01/14/98 - 0000Z (UTC) 01/15/98 Daily Status Report as of 015/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 7802 (Daily Darks to Update Acquisition Bad Pixel Table) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations to obtain daily darks to update the acquisition bad pixel table. These darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Seven Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed NIC/1/2/3 7318 (Ages of Extreme Metallicity Globular Clusters Near the Galactic Center) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the globular cluster NGC-104 which is located near our galactic center in order to determine the age of this extreme metallicity cluster. Failure to detect an age difference will imply that star formation and chemical enrichment occurred rapidly in the inner halo if the clusters shared a common origin. The observations were completed as planned, and no problems were reported. 1.4 Completed Two Sets of NIC/1/2/3 7456 (Water in Protoplanetary Disks) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the protoplanetary disk around the star RW-Aur to search for water in its disks. This survey will be the first to take a detailed look at the gas phase properties of pre-main-sequence disks at AU distances. The observations were completed as scheduled, and no anomalies were reported. 1.5 Completed Three Sets of NIC/1/3 7813 (NIC/3 HDF Darks) The Near Infrared Camera (Cameras One and Three) was used to make internal dark calibration images which will be used when processing the Hubble Deep Field images during the first NICMOS Camera Three campaign. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Two Sets of NIC/1/2/3 7322 (The IMF Below 1 Msun in Young Clusters: 1.9micron Water Band Imaging of IC348) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of cluster IC348. This study will be the first to obtain spectral types for a young cluster population complete to K=16, and thereby quantify, to unprecedentedly high precision, the IMF below 1M{sun} in young clusters. The observations were executed as scheduled, and no anomalies were reported. 1.7 Completed NIC/3 7816 (NIC/3 Photometric Zero Point) The Near Infrared Camera (Camera Three) was used to make observations of Standard star G191B2B-CA in order to study the zero point flux for a set of filters. The NIC/3 intra-pixel sensitivity will also be measured. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 07 Successful : 07 Scheduled Reacquisitions : 11 Successful : 11 2.2 FHST Updates: Scheduled : 07 Successful : 07 2.3 Operations Notes: No operational problems were reported and no HSTARs were written. Dumped SSR Interrupt Log at 014/1924Z. AEDP - Data processed through 014/2000Z and transmitted to PODPS at 015/0600Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH